OpenStudy (anonymous):

Which of these arcs has a measure of 134°? arc FJ arc DF arc EG arc DH

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

arc DF

OpenStudy (anonymous):

arc EF = 90 arc FG = 46 so arc DE = 180 -(90+46) arc DF = 90+ arc DE